OpenStudy (juan1857):

Write -2x+4y=9 in polar form using the polar form of a linear equation.

OpenStudy (juan1857):

i'm troubled, getting phi. I don't know when to add 180 to the Arctan

jimthompson5910 (jim_thompson5910):

You'll use these two equations \[\Large x = r\cos(\theta)\] \[\Large y = r\sin(\theta)\] \[\Large -2x + 4y = 9\] \[\Large -2*r\cos(\theta) + 4*r\sin(\theta) = 9\] Now solve for r

OpenStudy (juan1857):

ok so for p I got \[\frac{ -9\sqrt{5} }{ 10}\]

OpenStudy (juan1857):

and for phi, i got -63.43 @jim_thompson5910

OpenStudy (juan1857):

should I add 180, and can you explain to me when should I add it

jimthompson5910 (jim_thompson5910):

how are p and phi defined?

OpenStudy (juan1857):

no idea